OpenStudy (anonymous):

wat are the roots of 7x^2+392=0 can someone help will give medal

OpenStudy (solomonzelman):

no real solut

OpenStudy (freckles):

Can you isolate the x^2 part

OpenStudy (anonymous):

idk um so confused @freckles

OpenStudy (anonymous):

i dont know how to do it like this

OpenStudy (freckles):

\[ax^2+b=0 \\ \text{ subtracting } b \text{ on both sides gives } \\ ax^2+b-b=0-b \\ \text{ Since } b-b=0 \text{ and } 0-b=-b \\ \text{ then we have } ax^2+0=-b \\ \text{ Since } u+0=u \text{ then we have } \\ ax^2=-b \\ \text{ now to undo the operation of multiplication you divide both sides by something } \\ \text{ divide both sides by } a \\ \frac{ax^2}{a}=\frac{-b}{a} \\ x^2=\frac{-b}{a} \text{ since } \frac{a}{a}=1 \\ \]

OpenStudy (freckles):

take these exact steps to isolate your x^2

OpenStudy (freckles):

After you are done isolating just x^2 you can solve for x by square rooting both sides